# Is Your Paper Being Reviewed by an LLM? Benchmarking AI Text Detection in Peer Review

### v2

@@ 5.4 How do human-written and AI-generated peer reviews differ?
Prior work has shown that peer reviews written by GPT-4 and humans have a level of semantic similarity which is comparable to that between different human-authored peer reviews, which has been used to advocate for the usefulness of feedback from GPT-4 in the paper writing process [37]. In our qualitative analysis, we found that GPT-4 does indeed generate similar higher-level comments as human reviewers, which could account for this semantic similarity. Despite being generic in nature, we would agree that such feedback could be useful to authors seeking to improve their manuscripts. Nevertheless, we believe that the lack of specificity, detail, and consideration of related work in peer reviews authored by GPT-4 demonstrates that it is not suitable for replacing human domain experts in the peer review process.
@@ 5.5 Do AI-generated reviews assign higher scores than human reviews?
All LLMs produce higher scores than human reviews with a high degree of statistical significance, assessed using a two-sided Wilcoxon signed-rank test (see legend for $p$-values in Figure S3). While the difference between human and AI confidence scores are relatively consistent across all three LLMs, Claude exceeds human scores by the greatest magnitude for soundness, presentation, and contribution. GPT-4o and and Gemini exceed human scores by a similar magnitude for presentation and contribution, while GPT-4o exhibits a greater divergence for soundness scores. Overall these results indicate that AI-generated peer reviews are more favorable w.r.t. assigned scores than human-written peer reviews, which raises fairness concerns as scores are highly correlated with acceptance decisions. Our findings are consistent with prior work which has shown that papers reviewed by LLMs have a higher chance of acceptance [10, 6, 11].
6 Limitations
@@ 7 Conclusion
Appendix A Baseline methods.
We compare our approach to 18 baseline methods from IGMBT [^3] https://github.com/kinit-sk/IMGTB (released under MIT license) with its default setting [34], which are categorized into metric-based and pretrained model-based methods. The metric-based methods include Binoculars  [55], DetectLLM-LLR  [36], DNAGPT  [49], Entropy  [31], FastDetectGPT  [56], GLTR  [31], LLMDeviation  [57], Loglikelihood  [27], LogRank  [58], MFD  [57], Rank  [31], and S5  [34]. The model-based methods include NTNU-D  [59], ChatGPT-D  [20], OpenAI-D  [27], OpenAI-D-lrg  [27], RADAR-D  [27], and MAGE-D  [60].
@@ A.1.1 Binoculars
Binoculars [55] analyzes text through two perspectives. First, it calculates the log perplexity of the text using an observer LLM. Then, a performer LLM generates next-token predictions, whose perplexity is evaluated by the observer—this metric is termed cross-perplexity. The ratio of perplexity to cross-perplexity serves as an indicator for detecting LLM-generated text.
@@ A.1.2 DNAGPT
DNAGPT [49] is a training-free detection method designed to identify machine-generated text. Unlike conventional approaches that rely on training models, DNAGPT uses Divergent N-Gram Analysis (DNA) to detect discrepancies in text origin. The method works by truncating a given text at the midpoint and using the preceding portion as input to an LLM to regenerate the missing section. By comparing the regenerated text with the original through N-gram analysis (black-box) or probability divergence (white-box), DNAGPT reveals distributional differences between human and machine-written text, offering a flexible and explainable detection strategy.
@@ A.1.3 Entropy
Similar to the Rank score, the Entropy score for a text is determined by averaging the entropy values of each word, conditioned on its preceding context  [31].
@@ A.1.4 GLTR
The Entropy score, like the Rank score, is computed by averaging the entropy values of each word within a text, considering the preceding context  [31].
@@ A.1.5 MFD
The Multi-Feature Detection (MFD) method [57] detects AI-generated text using four features: log-likelihood, log-rank, entropy, and LLM deviation.
@@ A.1.6 Loglikelihood
This method utilizes a language model to compute the token-wise log probability. Specifically, given a text, the log probability of each token is averaged to produce a final score. A higher score indicates a greater likelihood that the text is machine-generated  [27].
@@ A.1.7 LogRank
Unlike the Rank metric, which relies on absolute rank values, the Log-Rank score is derived by applying a logarithmic function to the rank value of each word  [58].
@@ A.1.8 Rank
The Rank score is calculated by determining the absolute rank of each word in a text based on its preceding context. The final score is obtained by averaging the rank values across the text. A lower score suggests a higher probability that the text was machine-generated  [31].
@@ A.1.9 DetectLLM-LLR
This approach integrates Log-Likelihood and Log-Rank scores, leveraging their complementary properties to analyze a given text  [36].
@@ A.1.10 FastDetectGPT
This method assesses changes in a model’s log probability function when small perturbations are introduced to a text. The underlying idea is that LLM-generated text often resides in a local optimum of the model’s probability function. Consequently, minor perturbations to machine-generated text typically result in lower log probabilities, whereas perturbations to human-written text may lead to either an increase or decrease in log probability  [58].
@@ A.2.1 NTNU-D
It is a fine-tuned classification model based on the RoBERTa-base model, and three sizes of the bloomz-models  [59]
@@ A.2.2 ChatGPT-D
The ChatGPT Detector  [20] is designed to differentiate between human-written text and content generated by ChatGPT. It is based on a RoBERTa model that has been fine-tuned for this specific task. The authors propose two training approaches: one that trains the model solely on generated responses and another that incorporates both question-answer pairs for joint training. In our evaluation, we adopt the first approach to maintain consistency with other detection methods.
@@ A.2.3 OpenAI-D and RADAR-D
OpenAI Detector [27] is a fine-tuned RoBERTa model designed to identify outputs generated by GPT-2. Specifically, it was trained using text generated by the largest GPT-2 model (1.5B parameters) and is capable of determining whether a given text is machine-generated.
@@ A.2.4 MAGE-D
MAGE (MAchine-GEnerated text detection)  [60] is a large-scale benchmark designed for detecting AI-generated text. It compiles human-written content from seven diverse writing tasks, including story generation, news writing, and scientific writing. Corresponding machine-generated texts are produced using 27 different LLMs, such as ChatGPT, LLaMA, and Bloom, across three representative prompt types.
@@ B.1 Dataset
Our dataset is hosted on Hugging Face and publicly available at https://huggingface.co/datasets/IntelLabs/AI-Peer-Review-Detection-Benchmark. It is released under the Intel OBL Internal R&D Use License Agreement (https://huggingface.co/datasets/IntelLabs/AI-Peer-Review-Detection-Benchmark/blob/main/LICENSE.md).
@@ D.1 Voting Mechanism for the Anchor Embedding Method
D.2 Assessing Dataset Robustness to Prompt Variation (Supplementary)
@@
D.3 Examples from human analysis of differences between human and AI-written peer reviews
@@
D.4 Comparison of numeric scores assigned by human and AI reviewers
@@
Figure S3: Difference between AI and human scores. For each matched review (aligned by paper ID and recommendation), score differences were computed and displayed as histograms. Scores range from 1 to 4 for all metrics except Confidence, which ranges from 1 to 5. Statistical significance was assessed using a two-sided Wilcoxon signed-rank test, with p-values shown in the legend. This figure includes only NeurIPS2022–2024 and ICLR2024, because they are the onyl conferences that required reviewers to submit these scores in their review templates.
D.5 Validation of AI-edited reviews

### v3

@@ 5.4 How do human-written and AI-generated peer reviews differ?
Prior work has shown that peer reviews written by GPT-4 and humans have a level of semantic similarity which is comparable to that between different human-authored peer reviews, which has been used to advocate for the usefulness of feedback from GPT-4 in the paper writing process (Liang et al., 2024b). In our qualitative analysis, we found that GPT-4 does indeed generate similar higher-level comments as human reviewers, which could account for this semantic similarity. Despite being generic in nature, we would agree that such feedback could be useful to authors seeking to improve their manuscripts. Nevertheless, we believe that the lack of specificity, detail, and consideration of related work in peer reviews authored by GPT-4 demonstrates that it is not suitable for replacing human domain experts in the peer review process.
A natural concern is that the deficiencies of AI peer reviews (e.g., lack of references, generic feedback) could be artifacts of prompting, potentially affecting detectability. We tested this by generating 6,000 reviews with prompts updated to incorporate our findings above, explicitly instructing models to avoid typical AI review characteristics and to mimic more human-like reviews (Appendix E.5). Detection performance was essentially unchanged (AUROC differences $\leq$ 0.01), suggesting that (1) the gap between human and AI reviews is systematic rather than prompt-dependent, and (2) our results are consistent with the prompt-robustness analysis in Section 5.1.
@@ 5.5 Do AI-generated reviews assign higher scores than human reviews?
All LLMs produce higher scores than human reviews with a high degree of statistical significance, assessed using a two-sided Wilcoxon signed‐rank test (see legend for $p$-values in Figure S3). While the difference between human and AI confidence scores are relatively consistent across all three LLMs, Claude exceeds human scores by the greatest magnitude for soundness, presentation, and contribution. GPT-4o and and Gemini exceed human scores by a similar magnitude for presentation and contribution, while GPT-4o exhibits a greater divergence for soundness scores. Overall these results indicate that AI-generated peer reviews are more favorable w.r.t. assigned scores than human-written peer reviews, which raises fairness concerns as scores are highly correlated with acceptance decisions. Our findings are consistent with prior work which has shown that papers reviewed by LLMs have a higher chance of acceptance (Drori & Te’eni, 2024; Latona et al., 2024; Ye et al., 2024).
@@ 6 Conclusion
Appendix A Baseline methods
We compare our approach to 18 baseline methods from IGMBT [^3] https://github.com/kinit-sk/IMGTB (released under MIT license) with its default setting (Spiegel & Macko, 2023), which are categorized into metric-based and pretrained model-based methods. The metric-based methods include Binoculars  Hans et al. (2024), DetectLLM-LLR  Su et al. , DNAGPT  Yang et al. (2023), Entropy  Gehrmann et al. (2019), FastDetectGPT  Bao et al. (2023), GLTR  Gehrmann et al. (2019), LLMDeviation  Wu & Xiang (2023), Loglikelihood  Solaiman et al. (2019), LogRank  Mitchell et al. (2023a), MFD  Wu & Xiang (2023), Rank  Gehrmann et al. (2019), and S5  Spiegel & Macko (2023). The model-based methods include NTNU-D  Sivesind & Winje (2023), ChatGPT-D  Guo et al. (2023), OpenAI-D  Solaiman et al. (2019), OpenAI-D-lrg  Solaiman et al. (2019), RADAR-D  Solaiman et al. (2019), and MAGE-D  Li et al. (2024).
@@ A.1.1 Binoculars
Binoculars Hans et al. (2024) analyzes text through two perspectives. First, it calculates the log perplexity of the text using an observer LLM. Then, a performer LLM generates next-token predictions, whose perplexity is evaluated by the observer—this metric is termed cross-perplexity. The ratio of perplexity to cross-perplexity serves as an indicator for detecting LLM-generated text.
@@ A.1.2 DNAGPT
DNAGPT Yang et al. (2023) is a training-free detection method designed to identify machine-generated text. Unlike conventional approaches that rely on training models, DNAGPT uses Divergent N-Gram Analysis (DNA) to detect discrepancies in text origin. The method works by truncating a given text at the midpoint and using the preceding portion as input to an LLM to regenerate the missing section. By comparing the regenerated text with the original through N-gram analysis (black-box) or probability divergence (white-box), DNAGPT reveals distributional differences between human and machine-written text, offering a flexible and explainable detection strategy.
@@ A.1.3 Entropy
Similar to the Rank score, the Entropy score for a text is determined by averaging the entropy values of each word, conditioned on its preceding context  Gehrmann et al. (2019).
@@ A.1.4 GLTR
The Entropy score, like the Rank score, is computed by averaging the entropy values of each word within a text, considering the preceding context  Gehrmann et al. (2019).
@@ A.1.5 MFD
The Multi-Feature Detection (MFD) method Wu & Xiang (2023) detects AI-generated text using four features: log-likelihood, log-rank, entropy, and LLM deviation.
@@ A.1.6 Loglikelihood
This method utilizes a language model to compute the token-wise log probability. Specifically, given a text, the log probability of each token is averaged to produce a final score. A higher score indicates a greater likelihood that the text is machine-generated  Solaiman et al. (2019).
@@ A.1.7 LogRank
Unlike the Rank metric, which relies on absolute rank values, the Log-Rank score is derived by applying a logarithmic function to the rank value of each word  Mitchell et al. (2023a).
@@ A.1.8 Rank
The Rank score is calculated by determining the absolute rank of each word in a text based on its preceding context. The final score is obtained by averaging the rank values across the text. A lower score suggests a higher probability that the text was machine-generated  Gehrmann et al. (2019).
@@ A.1.9 DetectLLM-LLR
This approach integrates Log-Likelihood and Log-Rank scores, leveraging their complementary properties to analyze a given text  Su et al. .
@@ A.1.10 FastDetectGPT
This method assesses changes in a model’s log probability function when small perturbations are introduced to a text. The underlying idea is that LLM-generated text often resides in a local optimum of the model’s probability function. Consequently, minor perturbations to machine-generated text typically result in lower log probabilities, whereas perturbations to human-written text may lead to either an increase or decrease in log probability  Mitchell et al. (2023a).
@@ A.2.1 NTNU-D
It is a fine-tuned classification model based on the RoBERTa-base model, and three sizes of the bloomz-models  Sivesind & Winje (2023)
@@ A.2.2 ChatGPT-D
The ChatGPT Detector  Guo et al. (2023) is designed to differentiate between human-written text and content generated by ChatGPT. It is based on a RoBERTa model that has been fine-tuned for this specific task. The authors propose two training approaches: one that trains the model solely on generated responses and another that incorporates both question-answer pairs for joint training. In our evaluation, we adopt the first approach to maintain consistency with other detection methods.
@@ A.2.3 OpenAI-D and RADAR-D
OpenAI Detector Solaiman et al. (2019) is a fine-tuned RoBERTa model designed to identify outputs generated by GPT-2. Specifically, it was trained using text generated by the largest GPT-2 model (1.5B parameters) and is capable of determining whether a given text is machine-generated.
@@ A.2.4 MAGE-D
MAGE (MAchine-GEnerated text detection)  Li et al. (2024) is a large-scale benchmark designed for detecting AI-generated text. It compiles human-written content from seven diverse writing tasks, including story generation, news writing, and scientific writing. Corresponding machine-generated texts are produced using 27 different LLMs, such as ChatGPT, LLaMA, and Bloom, across three representative prompt types.
@@ B.1 Dataset
A subset of our dataset is included as part of the supplementary materials in a zipped file. The full dataset will be released publicly upon publication under a permissive research-use license.
@@ D.1 Voting Mechanism for the Anchor Embedding Method
D.2 Assessing Dataset Robustness to Prompt Variation
@@
D.3 Assessing Dataset Robustness To Agent-Based and Structured Generation Pipelines
To assess whether AI-generated reviews produced via more complex prompting pipelines differ in detectability, we evaluated Binoculars on reviews generated by recent agent-based and structured LLM systems. These include:
  • AI Scientist (Lu et al., 2024): Combines self-reflection, ensembling, and few-shot prompting to generate reviews with both simple and complex reasoning flows.
  • AgentReview (Jin et al., 2024): Uses GPT-4 agents simulating different reviewer archetypes (e.g., knowledgeable, irresponsible) in multi-turn interactions.
  • DeepReview (Zhu et al., 2025): Uses a structured, multi-stage pipeline with separate modules for scoring, justification, and recommendations.
These datasets only contain LLM-generated reviews. To compare them against human-written reviews, we paired them with our ICLR 2022 human review samples. Due to resource constraints and to minimize experimental overhead, we focus this test on Binoculars, the top-performing baseline amongst the existing AI text detection methods. As a control, we included samples from our own GPT-4o-generated reviews (ICLR 2022).
Table S11: AUROC of Binoculars on reviews from agent-based and structured LLM pipelines, paired with ICLR 2022 human-written reviews. The control row includes GPT-4o-generated reviews from our dataset for the same year. Variants of each review generation pipeline are indicated in parentheses. Mean, minimum, and maximum values are computed via bootstrapping (n=1,000).
Source Mean AUROC Min Max
AI Scientist (simple) 0.9955 0.9936 0.9975
AI Scientist (complex) 0.9973 0.9933 0.9992
AgentReview (baseline) 0.9953 0.9935 0.9964
AgentReview (benign) 0.9984 0.9974 0.9990
AgentReview (responsible) 0.9983 0.9976 0.9991
AgentReview (irresponsible) 0.9984 0.9975 0.9991
AgentReview (knowledgeable) 0.9982 0.9974 0.9990
AgentReview (unknowledgeable) 0.9985 0.9977 0.9991
DeepReview (fast) 0.9937 0.9918 0.9954
DeepReview (best) 0.9929 0.9914 0.9950
Control (GPT-4o) 0.9647 0.9604 0.9686
Across all agentic and structured generation pipelines tested, Binoculars consistently achieved high AUROC — often outperforming its results on our own GPT-4o samples (Table S11). This suggests that the detectability of AI-generated peer reviews is not fundamentally diminished by the use of agent-based workflows, CoT prompting, or structured review pipelines. While we limit this evaluation to one method due to computational constraints, the results strengthen the conclusion that our benchmark remains robust and relevant, even as more complex AI reviewing systems emerge.
D.4 Examples from human analysis of differences between human and AI-written peer reviews
@@
D.5 Prompt Enrichment Experiment
In Section 5.4, we identified several characteristic deficiencies of AI peer reviews compared to human-written ones, such as a lack of references to specific details in the paper, limited engagement with related work, and reliance on generic or boilerplate phrasing. One natural question is whether whether these deficiencies could be artifacts of the original prompting strategy, and whether updating the prompts to explicitly address them might close the gap with human-authored reviews. To test this, we added an extra prompt to the original one that incorporated our findings from Section 5.4, instructing models to avoid these AI-like characteristics and to produce reviews that more closely resemble human ones. The following is the added prompt:
Do the following: {internallinenumbers*} - Reference specific elements from the paper (e.g., particular sections, tables, figures, equations, or experimental results). {internallinenumbers*} - Engage with related work: mention missing citations, suggest relevant prior studies, and evaluate the paper’s novelty in the context of existing literature. {internallinenumbers*} - Provide paper-specific, varied feedback rather than relying on template-like language. - Balance high-level assessments with detailed, grounded observations. Do NOT do the following: - Do not give only vague or generic comments without specific evidence from the paper. - Do not recycle identical phrasing or boilerplate critiques across reviews. - Do not omit discussion of related work or novelty considerations. {internallinenumbers*} - Do not provide feedback that could apply equally to almost any paper; make each review unique and tailored.
Using this revised prompt, we generated 1,000 reviews for papers in our calibration set across four conference-year pairs (ICLR and NeurIPS 2021–2022) using the same three commercial LLMs from our study (GPT‑4o, Gemini, Claude). With matching human reviews, this produced a total of 6,000 reviews. We then compared the detection performance of the top two performing methods (Anchor and Binoculars) between the original and enriched prompt settings.
Detector Original Prompt Enriched Prompt
Anchor 0.995 0.997
Binoculars 0.982 0.977
Table S13: Detection performance (AUROC) under original and enriched prompting strategies.
As shown in Table S13, detection performance remained virtually unchanged (AUROC differences within 0.01), indicating that (1) the distinction between human- and AI-written reviews reflects a systematic gap rather than a prompt artifact, and (2) the outcome aligns with the prompt-robustness analysis in Section 5.1.
D.6 Comparison of numeric scores assigned by human and AI reviewers
@@
Figure S3: Difference between AI and human scores. For each matched review (aligned by paper ID and recommendation), score differences were computed and displayed as histograms. Scores range from 1 to 4 for all metrics except Confidence, which ranges from 1 to 5. Statistical significance was assessed using a two-sided Wilcoxon signed‐rank test, with p-values shown in the legend. This figure includes only NeurIPS2022–2024 and ICLR2024, because they are the onyl conferences that required reviewers to submit these scores in their review templates.
D.7 Validation of AI-edited reviews
@@
D.8 Sensitivity of Anchor embedding detectability performance to embedding model
In addition to the OpenAI embedding model used for our main Anchor embedding results, we conducted experiments using two other embeddings models: one based on a traditional encoder architecture (Jina AI embedding model v3 (Sturua et al., 2024)), and another based on a decoder architecture (Intfloat e5-mistral-7b-instruct (Wang et al., 2023; 2022)). We picked these based on their popularity (assessed via hugging face downloads) and their ability to accommodate long context lengths. The evaluation results for the Anchor embedding method when used with these models on the test set are provided in Table S15. While we observe that the performance of the anchor embedding approach is lower using these alternative embedding models than when utilizing the OpenAI embedding model, its performance is still comparable to that of Binoculars. These results illustrate how the choice of the embedding model can impact detection performance of the Anchor approach.
Table S15: Anchor method detectability results using other embedding models
Embedding Model Test Reviews Target FPR Actual FPR Actual TPR
Jina GPT-4o 0.1% 0.2% 50.5%
Jina GPT-4o 0.5% 0.7% 64.8%
Jina GPT-4o 1.0% 1.1% 71.4%
Jina Gemini 0.1% 0.2% 29.7%
Jina Gemini 0.5% 0.7% 46.2%
Jina Gemini 1.0% 1.1% 53.1%
Jina Claude 0.1% 0.2% 22.0%
Jina Claude 0.5% 0.7% 36.9%
Jina Claude 1.0% 1.1% 45.1%
e5-mistral GPT-4o 0.1% 0.1% 26.0%
e5-mistral GPT-4o 0.5% 0.4% 41.7%
e5-mistral GPT-4o 1.0% 0.8% 48.4%
e5-mistral Gemini 0.1% 0.1% 26.4%
e5-mistral Gemini 0.5% 0.4% 40.8%
e5-mistral Gemini 1.0% 0.8% 47.4%
e5-mistral Claude 0.1% 0.1% 26.6%
e5-mistral Claude 0.5% 0.4% 43.9%
e5-mistral Claude 1.0% 0.8% 51.9%
D.9 Impact of finetuning supervised classifiers on AI peer review dataset
Our main results in Table 2 showed that existing detection methods which are based on finetuned classifiers perform poorly at detecting AI text in peer reviews. One reason for this could be due to a mismatch between the domain of text used to train these classifiers and our task of detecting AI text in the peer review domain. Therefore, we conducted experiments where we finetuned supervised classifiers on our dataset of paired human & AI-written peer reviews.
As the existing methods based supervised classifiers are all finetuned variants of RoBERTa, we similarly trained a RoBERTa classifier separately on peer reviews generated by GPT-4o, Claude, and Gemini to assess their out-of-domain detection accuracy. To maintain a fair comparison, we used the exact same calibration dataset (consisting of GPT-4o, Claude, and Gemini reviews) to calibrate the classification threshold as was used for our anchor embedding approach. The performance of the resulting classifiers is summarized in Table S16.
Table S16: Performance of RoBERTa classifiers trained & tested on different subsets of AI-generated peer reviews from our dataset
Train Reviews Test Reviews Target FPR Actual FPR Actual TPR
GPT-4o GPT-4o 0.1% 0.3% 100%
GPT-4o GPT-4o 0.5% 0.6% 100%
GPT-4o GPT-4o 1.0% 0.9% 100%
GPT-4o Gemini 0.1% 0.3% 83.5%
GPT-4o Gemini 0.5% 0.6% 92.0%
GPT-4o Gemini 1.0% 0.9% 95.8%
GPT-4o Claude 0.1% 0.3% 44.9%
GPT-4o Claude 0.5% 0.6% 59.1%
GPT-4o Claude 1.0% 0.9% 70.4%
Gemini GPT-4o 0.1% 0.4% 99.9%
Gemini GPT-4o 0.5% 0.8% 99.6%
Gemini GPT-4o 1.0% 1.4% 99.9%
Gemini Gemini 0.1% 0.4% 100%
Gemini Gemini 0.5% 0.8% 100%
Gemini Gemini 1.0% 1.4% 100%
Gemini Claude 0.1% 0.4% 41.7%
Gemini Claude 0.5% 0.8% 49.8%
Gemini Claude 1.0% 1.4% 60.2%
Claude GPT-4o 0.1% 0.5% 95.9%
Claude GPT-4o 0.5% 0.7% 97.6%
Claude GPT-4o 1.0% 0.9% 98.3%
Claude Gemini 0.1% 0.5% 2.7%
Claude Gemini 0.5% 0.7% 3.0%
Claude Gemini 1.0% 0.9% 3.5%
Claude Claude 0.1% 0.5% 100%
Claude Claude 0.5% 0.7% 100%
Claude Claude 1.0% 0.9% 100%
When tested on AI reviews generated by the same LLM as was used to construct the training dataset, we observe that these classifiers achieve near 100% TPR at an FPR < 1%. However, when evaluated on out-of-domain test data (i.e., peer reviews generated by LLMs other than the one used to construct the training dataset), the performance is highly variable and inconsistent. Thus, the performance of supervised methods is dependent upon the training dataset including samples from the same LLM which was used to generate the evaluated text.
Using the same calibration dataset as these supervised finetuning experiments, our Anchor embedding approach achieves overall strong performance across reviews generated by all LLMs without requiring any separate training. We also note that training supervised classifiers to achieve competitive performance in this task would not be possible without our paired human & AI-written peer review dataset, which is a core contribution of our work. Therefore, the improved performance of supervised methods when trained on our dataset (relative to their baseline implementations trained on general texts) can be viewed as further evidence of our dataset’s value.
D.10 Topic Distribution of Papers for Evaluated Reviews
We analyzed the keywords associated with 500 papers from our ICLR 2021 test set using metadata available for each paper in OpenReview. The statistic of the topics is given below.
  1. 1.
    Uncategorized: 19.5%
  2. 2.
    Deep Learning Foundations: 17.5%
  3. 3.
    Reinforcement Learning: 12.9%
  4. 4.
    Robustness & Adversarial M: 11.2%
  5. 5.
    Representation Learning: 8.0%
  6. 6.
    Optimization & Training: 8.0%
  7. 7.
    NLP & Multimodal: 6.0%
  8. 8.
    Computer Vision: 5.7%
  9. 9.
    Meta-Learning: 4.6%
  10. 10.
    Theory & Interpretability: 3.7%
  11. 11.
    Model Efficiency: 2.6%
This breakdown indicates that our benchmark is not dominated by any single domain, and therefore the observed differences between human-written and LLM-generated reviews cannot be attributed to a narrow topical bias in the underlying papers.
D.11 How do common factors which are characteristic of AI peer reviews impact detectability?
Building upon our analysis of how human and AI-written peer reviews differ (Section 5.4 and Appendix D.4), we analyzed AI-written reviews from the ICLR 2021 test set for the presence of three common factors that we previously identified as characteristic of AI reviews: (1) a lack of references to prior work, (2) a lack of references to specific details in the paper, and (3) overly generic criticisms. We used an LLM-as-a-judge approach with GPT-4.1 to quantify which of these issues were present across the 500 analyzed reviews.
The most commonly identified factor among this set of reviews was the lack of references to prior work, which was present in over 99% of the cases where the review was flagged as AI-written by the Anchor method. The next most common factor was the lack of references to specific details in the paper, which was flagged in 26% of papers which were detected as AI-written. Finally, 22% of papers classified as AI-generated were determined to have overly generic criticisms. These results indicate that a lack of reference to prior work is the most prevalent factor present among AI-generated reviews.
Interestingly, we do not observe a correlation between the presence of these factors in an AI-generated review and the detection accuracy of the Anchor embedding method. This could be due to the already strong detection accuracy of the Anchor method which limits the number of false negatives (only 14.4% of AI-generated reviews were classified as human-written at a calibrated FPR of 0.1%). An alternative explanation could be that the Anchor embedding method relies on other, more nuanced linguistic differences than the factors which humans identify as being characteristic of AI-generated reviews (e.g., lack of references, generic criticisms).